You are analyzing a project which has a $20,000 sunk cost. The project’s cost is $150,000 and the expected cash flows are: (Use this info for questions #1-4)

            Year 1: $60,000

            Year 2: $0

            Year 3: $100,000

            Year 4: $0

            Year 5: $60,000

The discount rate is 18% (the required return is also 18%). The required payback is 2.3 years.

1. Calculate the internal rate of return (must have correct set-up and your answer should be between two consecutive whole percentages). Accept or reject?

2. Calculate the discounted payback period. Accept or reject?

3. Calculate the payback period. Accept or reject?

4. Calculate the NPV.  Accept or reject?
Answer #1

1) IRR is the rate at which NPV is 0,Let us assume rate to be 14%

Statement showing NPV at 14%

Year Cash flow PVIF @ 14% Present value
1 60000 0.8772 52632
2 0 0.7695 0
3 100000 0.6750 67497
4 0 0.5921 0
5 60000 0.5194 31162
Total 151291
Less: Initial Investment 150000
NPV 1291

Now let us assume rate to be 15%,

Statement showing Npv at 15%

Year Cash flow PVIF @ 15% Present value
1 60000 0.8696 52174
2 0 0.7561 0
3 100000 0.6575 65752
4 0 0.5718 0
5 60000 0.4972 29831
Total 147756
Less: Initial Investment 150000
NPV -2244

Using interpolation one can find IRR

Rate NPV
14% 1291
15% -2244
1% up -3535
? -1291

=1291/3535 = 0.36

Thus IRR = 14+0.36

=14.36%

Since IRR is less than NPV, project should be rejected

2) Statement showing discounted cash payback period

Year Cash flow PVIF @ 18% Present value Cummulative Present value
1 60000 0.8475 50847.46              50,847.46
2 0 0.7182 0.00              50,847.46
3 100000 0.6086 60863.09            111,710.54
4 0 0.5158 0.00            111,710.54
5 60000 0.4371 26226.55            137,937.10

Total cummulative present value at end of year 5 is 137,937.1$ which is less that projects initial cost i.e 150,000$

hence project won't pay back even at end of year 5

one should reject the project

3) Statement showing payback period

Year Cash flow CummulatiCash flow
1 60000 60000
2 0 60000
3 100000 160000
4 0 160000
5 60000 220000

Using interpolation one can find payback period

Year Cummulative cash flow
2 60000
3 160000
1 up 100000
? 90000

=1*90000/1000000

=0.9

Thus payback period = 2 + 0.9 =2.9 years

Reject the project

4) Statement showing NPV

Year Cash flow PVIF @ 18% Present value
1 60000 0.8475 50847
2 0 0.7182 0
3 100000 0.6086 60863
4 0 0.5158 0
5 60000 0.4371 26227
Total 137937
Less: Initial Investment 150000
NPV -12063

Since NPV is negative , project should be rejected